On one day Fred's sport world sold 9 Pittsburgh steelers jerseys and 3 Cleveland Browns jerseys for a total of $899.40. The next day they sold 12 Pittsburgh steelers jerseys and 2 Cleveland Browns jerseys for a total of $1139.30. How much is the Pittsburgh steelers jersey and the Cleveland Browns jersey?
Answer by TimothyLamb(4379)

x = dollar price of steelers jersey
y = dollar price of browns jersey
---
9x + 3y = 899.40
12x + 2y = 1139.30
---
put the system of linear equations into standard form
---
9x + 3y = 899.40
12x + 2y = 1139.30
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = dollar price of steelers jersey = 89.95
y = dollar price of browns jersey = 29.95
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---